OpenStudy (anonymous):

Factoring Polynomials! Help Please! Jim You Too! Just 1 Question! Please!

OpenStudy (anonymous):

8r^3-64r^2+r-8=0 8r^2(r-8)+1(r-8)=0 (8r^2+1)(r-8)=0 I'll let you finish.

OpenStudy (anonymous):

Someone explain this much but i'm not understanding were he got the "+1". Explain?

OpenStudy (anonymous):

\[8r^{2} - 64r ^{2} + r - 8 = 0\]